OpenStudy (anonymous):

Find the domain and range of the function: f(x)= x^2-10x+7

OpenStudy (zehanz):

Domain: all the numbers you can put into f. Now ask yourself this: are there numbers that: - can't be squared? - can't be multiplied with -10 and then added 7? No! so there are no exceptions, the domain is R (all real numbers). The range is harder. f is a 2nd deg. polynomial, so the graph is a parabola. The vertex of the parabola is the highest or lowest point, so that sets a boundary for the range. You will have to find the vertex to get any further with this one. Do you know how to do that?

OpenStudy (anonymous):

-b/2a?

OpenStudy (zehanz):

That will do it, you now have to think about whether there is a max or min value for x = -b/2

OpenStudy (anonymous):

i think it is a minimum

OpenStudy (zehanz):

Right, it is a minimum, value is f(-b/(2a))= f(5) = 25 - 50 + 7 = -18. So the range is \[[-18, \rightarrow)\](or some other notation that means the same ;)